OpenStudy (anonymous):

if x=sqrt(-3)/2 , find the value of sqrt(1+x)+sqrt(1-x)/sqrt(1+x)-sqrt(1-x)

OpenStudy (dumbcow):

cant take the sqrt of neg num is it supposed to deal with complex num

OpenStudy (anonymous):

oh sorry, its written as x=sqrt(3)/2

OpenStudy (dumbcow):

i made a mistake it simplifies to \[\frac{1+\sqrt{1-x^{2}}}{x}\] \[x=\frac{\sqrt{3}}{2}\] \[=\sqrt{3}\]

OpenStudy (anonymous):

could you please provide some info on how we get 1+sqrt(1-x^2)/x ?

OpenStudy (dumbcow):

multiply by conjugate on top and bottom then combine like terms

OpenStudy (dumbcow):

conjugate is sqrt(1+x) + sqrt(1-x)

OpenStudy (anonymous):

got till 1+x+1-x+2sqrt(1-x^2)/(1-x)+(1-x)